# Understanding Why Label Smoothing Degrades Selective Classification and How to Fix It

### v1

@@ 0.D.1 On the Assumption that Logits are Positive
Figure 19: $v^{\prime}_{\text{max}}$ given $v_{\text{max}}$ within bins of similar MSP for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set. Generally, $v^{\prime}_{\text{max}}$ increases as $v_{\text{max}}$ decreases, showing empirically how logit normalisation increases uncertainty for higher $v_{\text{max}}$. We calculate the mean±$\pm$±std. of $v^{\prime}_{\text{max}}$ in a 0.2-wide sliding window for samples with $v_{\text{max}}$ within mean±$\pm$2std of $v_{\text{max}}$ within the bin (to remove noisy averages).
@@
Result 1 assumes that all elements in the logit vector are $>0$, which is not necessarily true. However, we also find empirically that both $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ tend to be dominated by the largest positive logits. This is intuitive as exponentiating or raising to power $p$ ($p\geq 4$ in our experiments) will amplify the larger logits. This is shown in Fig. 18, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set. Thus $\pi_{\text{max}}(\boldsymbol{v})\approx\pi_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$ and $v^{\prime}_{\text{max}}(\boldsymbol{v})\approx v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, where the top-$k$ logits are positive. As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to penalise higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$.
@@
We can consider the scenario where we add $\eta$ to only the top-$k$ logits. This more aptly describes the empirical logit behaviour compared to adding $\eta$ to all logits, as the values of the lower ranking logits vary much less than the higher ranking ones (Fig. 18). Here we would expect $\pi_{\text{max}}(\boldsymbol{v})$ to increase very slightly, but would expect $v^{\prime}_{\text{max}}(\boldsymbol{v})$ to decrease as the numerator of Eq. 24 would be dominated by the top-$k$ largest logits.
@@
Fig. 19 shows how empirically logit normalisation indeed increases uncertainty for higher $v_{\text{max}}$. We plot the mean±$\pm$±std of $v^{\prime}_{\text{max}}$ given $v_{\text{max}}$ for samples in different MSP bins. We see clearly that in almost all cases, for samples with similar $\pi_{\text{max}}$, the normalised max logit $v^{\prime}_{\text{max}}$ increases as the original max logit $v_{\text{max}}$ decreases. As such, logit normalisation is able to improve the SC performance of LS-trained models.
0.D.2 Existing Benchmarks and Training Recipes
@@
Although we do not exhaustively search all training recipes for all models benchmarked in [20, 4], we do provide a number of examples of evaluated models trained with label smoothing. We also provide links to publicly available training repositories, as not all papers mention label smoothing even when it is used in training. Upon inspection of [20, 4], these models do in fact seem to underperform at selective classification (and Cattelan and Silva [4] report that their AURCs benefit from logit normalisation).
@@
Galil et al. [20] state that some of their best performing ViT models [13, 67, 8] are trained with label smoothing (their Tab.1). However, after inspecting both the original papers and open-source repositories [^6] https://github.com/google-research/vision_transformer of the aforementioned work we were unable to find any confirmation of the use of label smoothing.

### v2

@@ E.3 On the assumption that logits are positive
Figure 20: $v^{\prime}_{\text{max}}$ given $v_{\text{max}}$ within bins of similar MSP for ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set. Generally, $v^{\prime}_{\text{max}}$ decreases as $v_{\text{max}}$ increases, showing empirically how logit normalisation increases uncertainty for higher $v_{\text{max}}$. We calculate the mean±$\pm$±std. of $v^{\prime}_{\text{max}}$ in a 0.2-wide sliding window for samples with $v_{\text{max}}$ within mean±$\pm$2std of $v_{\text{max}}$ within the bin (to remove noisy averages).
@@
Result 1 assumes that all elements in the logit vector are $>0$, which is not necessarily true. However, we also find empirically that both $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ tend to be dominated by the largest positive logits. This is intuitive as exponentiating or raising to power $p>1$ will amplify the larger logits. This is shown in Fig. 19, where we plot the mean±$\pm$±std of $v,v^{5}$ and $\exp v$ for the sorted logits of ResNet-50 $\alpha=0.2$ on the ImageNet evaluation set ($p=5$ is optimal on the validation data for logit normalisation in this case). Thus $\pi_{\text{max}}(\boldsymbol{v})\approx\pi_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$ and $v^{\prime}_{\text{max}}(\boldsymbol{v})\approx v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, where the top-$k$ logits are positive. As Result 1 holds for $v^{\prime}_{\text{max}}(\boldsymbol{v}_{\text{top-}k})$, we still expect logit normalisation to increase uncertainty for higher $v_{\text{max}}$ when comparing samples with similar $\pi_{\text{max}}$.
@@
We can also consider the scenario where we add $\eta$ to only the top-$k$ logits. This more aptly describes the empirical logit behaviour compared to adding $\eta$ to all logits, as the values of the lower ranking logits vary much less than the higher ranking ones (Fig. 19). Here we would expect $\pi_{\text{max}}(\boldsymbol{v})$ to increase very slightly, but would expect $v^{\prime}_{\text{max}}(\boldsymbol{v})$ to decrease as the numerator of Eq. 26 would be dominated by the top-$k$ largest logits.
@@
Fig. 20 shows how empirically logit normalisation indeed increases uncertainty for higher $v_{\text{max}}$. We plot the mean±$\pm$±std of $v^{\prime}_{\text{max}}$ given $v_{\text{max}}$ for samples in different MSP bins. We see clearly that in almost all cases, for samples with similar $\pi_{\text{max}}$, the normalised max logit $v^{\prime}_{\text{max}}$ decreases as the original max logit $v_{\text{max}}$ increases. As shown in Figs. 5, 16, 11 and 12, LS leads to misclassifications ✗ having higher $v_{\text{max}}$ than correct predictions ✓. Thus, logit normalisation is able to improve the SC performance of LS-trained models by penalising the confidence of higher max logit values.
E.4 Existing benchmarks and training recipes with LS
@@
Although we do not exhaustively search all training recipes for all models benchmarked in (Galil et al., 2023; Cattelan & Silva, 2024), we do provide a number of examples of evaluated models trained with label smoothing. We also provide links to publicly available training repositories, as not all papers mention label smoothing even when it is used in training. Upon inspection of (Galil et al., 2023; Cattelan & Silva, 2024), these models do in fact seem to underperform at selective classification (and Cattelan & Silva (2024) report that their AURCs benefit from logit normalisation).
@@
Galil et al. (2023) state that some of their best performing (at SC) ViT models (Dosovitskiy et al., 2021; Steiner et al., 2022; Chen et al., 2022) are trained with label smoothing (their Tab.1). However, after inspecting both the original papers and open-source repositories [^7] https://github.com/google-research/vision_transformer of the aforementioned work we were unable to find any confirmation of the use of label smoothing.
E.5 Logit normalisation in (Cattelan & Silva, 2024)
Explanations of effectiveness.
The work that introduces logit normalisation (Cattelan & Silva, 2024) is primarily an experimental study, where the focus is on extracting empirical takeaways. However, the authors do discuss potential reasons for the effectiveness of the approach in their Appendix B. They observe that models that are generally (on both ✓ and ✗) more uncertain benefit from logit normalisation and so suggest that logit normalisation alleviates “underconfidence” by reducing the influence of lower-ranked logits on the uncertainty of a prediction (using analysis similar to ours in Appendix D). We believe their explanation is ultimately incomplete as:
  1. 1.
    They do not clearly delineate the definition of over/underconfidence used in model calibration with that used in selective classification. This is an issue since calibration is concerned with absolute marginal (averaged over data samples) properties, whilst selective classification is concerned with relative conditional (per sample) properties. It is possible to be extremely over/underconfident in the calibration sense, whilst being optimal for SC, or very well calibrated and worse at distinguishing correct vs incorrect samples Zhu et al. (2024).
  2. 2.
    Although they elucidate some of the mechanics of logit normalisation, showing that logit normalisation reduces the impact of smaller logits on uncertainty, they do not link it to any observed model behaviour that differs between correct ✓ vs incorrect ✗ predictions, instead just speculating that models may be underconfident on certain sets of samples. To explain why an approach improves SC, we need to know how it treats ✓ and ✗ differently to improve the rank ordering of uncertainties between ✓ and ✗.
On the other hand, our explanation focuses solely on the relative ranking of uncertainties between ✓ vs ✗ samples and is able to link the mechanics of logit normalisation to how the behaviour of correct ✓ vs incorrect ✗ predictions differ under label smoothing (Figs. 5 and 1).
Logit centralisation.
Cattelan & Silva (2024) suggest an additional step of logit centralisation,
$\boldsymbol{v}^{\prime}=\frac{\boldsymbol{v}-\mu(\boldsymbol{v})}{\left\lVert\boldsymbol{v}-\mu(\boldsymbol{v})\right\rVert_{p}},\quad\mu(\boldsymbol{v})=\frac{1}{K}\sum_{k}v_{k},$ (28)
which in practice brings most of the lower valued logits close to zero if they are not already. They note that for the majority of models this step is unnecessary, and we also find it unnecessary in our experiments so we omit it for the sake of simplicity. We note that according to the analysis in Sec. 5.1 and Sec. E.3, for logit normalisation to be effective, $v^{\prime}_{\text{max}}$ and $\pi_{\text{max}}$ should be dominated by positive logits, which centralisation seems to empirically ensure by making most of the smallest logits close to zero (Cattelan & Silva, 2024).
Appendix F Additional segmentation results
In this section, we complete the picture by providing additional semantic segmentation results to Figures 1 and 8. As in the main paper, the following predictions are performed by a ResNet-101-based DeepLab-v3+ trained on Cityscapes. In these figures, we provide the rank of all the predictions – not only errors – to show the difference between the model’s behavior on its errors but also correct predictions. The rank of the correct predictions ✓ is slightly greyed out compared to the rank of the errors ✗. Figure 21 presents the full ranks corresponding to the scene used in the main paper. We will provide other examples to the readers upon request. Later, we will make our models available to ease the reproducibility of our work.
Figure 21: We provide all the predicted ranks on our segmentation plots to understand where the model predicted higher relative confidence. The pixel uncertainty rankings provided by the model trained with CE (left) are smoother than those of the model trained with LS $\alpha=0.2$ (right).
Figure 22 shows that the cross-entropy-based model predicts label maps in which pixel uncertainty ranks are smoother than those of the label-smoothing-based one. The label-smoothing model exhibits large areas with low uncertainty rankings, which suddenly decreases nearby the boundary of objects. However, the boundaries are sometimes wrongly predicted, leading to (rank-wise) highly confident errors ✗. The pixels corresponding to the boundaries of objects, where naturally the probability of error is higher and which are most sensitive to label errors, have higher confidence (lower uncertainty ranks) according to the LS-based model than the CE-based model. This explains the high-confidence errors ✗ in the misclassified boundaries of several objects, either in Figs. 1 and 22, and aligns with the imbalanced regularisation of Eq. 15.
We also provide Figure 23 to show that the logit normalisation method is not always successful in solving the ranking of the predictions. In this figure, we see that a large part of the misclassified pixels (lower left) keeps low uncertainty ranks either using the CE-based model or the LS-based model while using logit normalisation. However, there are some improvements on the label-smoothing side, where the original low-ranked boundary errors ✗ are less (rankwise) confident, such as the roof of the car and the right border of the median strip. Interestingly, LS-based high confidence errors in Figure 23 mainly correspond to a change of texture of the median strip – likely associated with an uncertainty on the true label – and correlate with those of the CE-based model. For brevity, we provide an image for a logit normalisation with norm dimension $p=5$, but the results do not differ much from the other values. As stated above, we will provide additional figures to readers upon request.
Figure 22: Logit normalisation improves on boundary-related very confident errors ✗, but not on the bottom right errors ✗.
Figure 23: Logit normalisation does not always fix selective classification when used on a model trained with a label-smoothing (right). However, boundary errors ✗ have lower (ranked) confidence, such as for the car’s roof.